Answer the following question with a clear explanation, showing any steps or processes used to reach the answer. Explain your process as though you are teaching the concept to a student who isn't familiar with Business Management.

Activity Time Immediate Activityin Days Predecessor(s) 卡 多 4 15 9 9 多 C, D F, E

a. What is the critical path method and why is it useful when planning projects?
b. Determine the slack or float for each activity. What does slack mean?
c. Find the critical path and determine how long the project will take if there aren't any delays.

Pls write the answer with complete explanation
Answer #1

Ans: Critical Path is the longest path in the project network. A path is called critical when all the activities fall with in the network is performed before reaching to the end of the project. Critical path method is used to plan the project in efficient and effective. This helps in finding the duration of the project, allocation of resources and crashing of activity to finish the project within minimum time and cost. Diagram of the given question is given in the image attached

C-12 L-19 ヲ。 E 39 2 C- e 28 L- eo e- 3 L-3 L-13

b) Determining the Float or Slack

Activity

Duration

(Dij)

EST (Earliest start time)

(Ei)

LST (latest start time)

Li= (Lj-Dij)

EFT ( Earliest Finish time)

Ej=(Ei-Dij)

LFT ( Latest Finish time)

Lj

Total Float

(LST- EST)

Slack of head event

(Lj-Ej)

Slack of tail event

(Li- Ei)

Free Float

(Total Float-Slack of head event)  

A 3 0 0 3 3 0 0 0 0
B 6 3 3 9 9 0 0 0 0
C 8 9 11 17 19 2 2 0 0
D 4 9 9 13 13 0 0 0 0
E 15 13 13 28 28 0 0 0 0
F 9 17 19 26 28 2 0 2 2
G 9 28 28 37 37 0 0 0 0

Slack means the time gap between the Earliest start time with that of the Latest start time or Earliest finish time with that of the Latest finish time.

C) Critical path

A-B-D-E-G= 3+6+4+15+9 = 37

A-B-D-F-G= 3+6+4+9+9 = 31

A-B-C-E-G= 3+6+8+9+9 = 35

So Critical path is  A-B-D-E-G with duration of 37 days

Steps of finding critical path are

1) Draw the Network diagram.

2) Determine the Earliest start time, Earliest finish time, Latest start time and Latest finish time

3) Identify different paths in the network.

4) from identified paths in the network the path having the longest duration is the critical path